The number of people that enter Saint Bernard’s cathedral on any given day is a Poisson distributed random variable with mean 10. In the cathedral are 14 donation boxes that are emptied every night. Each time a visitor enters the cathedral, they place a donation in a randomly chosen donation box, with each box equally likely to be chosen. Find the expected number of empty donation boxes at the end of the day.
Answer #1

Answer:

Given,

p = 1/14

we know,

p + q = 1

q = 1 - p

= 1 - 1/14

= 13/14

sample n = 10

Expected number of empty donation boxes at the end of the day = nq

= 10*13/14

= 130/14

= 9.2857

= 9.3